OpenStudy (anonymous):

does anyone know why y'=2x, y(0)=0, y(1)=100 has no solution?

OpenStudy (jamesj):

If dy/dx = 2x, what is the general solution of that equation? It's y(x) = x^2 + C. Now is there ANY choice of the number C such that y(0) = 0 and y(1) = 100?

OpenStudy (anonymous):

Ya general solution y=x^2 + c1 no problem. but the book says "Show that the problem has no solution. Is this an Initial value problem?" I don't how this statement can be correct when c can take on any value since it is arbitrary.

myininaya (myininaya):

right you have a contradiction so there is no y with those conditions that satisfy the differential equation

OpenStudy (jamesj):

Given that y(x) = x^2 + c, you now what to find ONE value of c such that it both true that y(0) = 0 AND y(1) = 100 Does there exist one value of c such that both of those conditions are satisfied?

OpenStudy (anonymous):

nope

OpenStudy (jamesj):

Therefore the initial value problem y' = 2x, y(0) = 0, y(1) = 100 has no solution.

OpenStudy (anonymous):

why does it have to be so opaque, but yet so simple.

OpenStudy (jamesj):

it's not opaque now. That's why we do exercises! :-)